An even number is chosen randomly from a set of cards labeled with the
numbers 1 through 8. A second even number is chosen without the first
card being replaced. Are these independent or dependent events? What is
the probability of both events occurring?

Answer by rothauserc(4718)

Probability of first number chosen is even = 4/8 = 1/2
:
Probability of second number chosen is even without replacement = 3/7
:
These events are dependent, that is the outcome of one event affects the outcome of another
:
Probability of both chosen number are even = 1/2 * 3/7 = 3/14
: